Fred is 12 years older than his brother.
In 2 years Fred will be twice as old as his brother will be.
How old is each of them now?

Answer by ankor@dixie-net.com(22740)

Let f = Fred's age now
Let b = brother's age now
:
Write an equation for each statement:
:
"Fred is 12 years older than his brother."
f = b + 12
:
"In 2 years Fred will be twice as old as his brother will be."
f + 2 = 2(b+2)
f + 2 = 2b + 4
f = 2b + 4 - 2
f = 2b + 2
How old is each of them now?
:
Substitute (b+12) for f in the above equation:
b+12 = 2b + 2
b -2b = 2 - 12
-b = -10
b = 10 yrs is brothers age
then
12 + 10 = 22 yrs is Fred's age
;
:
Check solution in statement:
"In 2 years Fred will be twice as old as his brother will be.
22 + 2 = 2(10 +2)